# A Sizable Discrepancy in Ground-Based JAGB Distances to Nearby Galaxies

### v1

@@ 2 Comparison of Distance Scales
The comparison between the two samples is shown in the top-half of Figure 1. The plotted uncertainties in the JAGB distances are re-tabulated from Table 3 in Lee et al. (2024c) to include contributions from the 2MASS photometric zeropoint and 2MASS-to-FourStar correction, error on the JAGB mode, choice of smoothing scale, and a 15 uncertainty in the foreground reddening (). We also include JAGB results for WLM (Lee et al., 2021) and M33 (Lee et al., 2022), which are part of the same ground-based observing program, for a total of thirteen galaxies for comparison. The uncertainties in the TRGB distances are re-tabulated from the CMDs/TRGB catalog to include the statistical uncertainty in the tip measurement and the same 15 uncertainty in foreground reddening (). The mean difference between the two sets of distances is = 0.17 0.04 (stat) 0.06 (sys) mag. The statistical uncertainty is given by , and the systematic uncertainty is given by the quadrature sum of the underlying calibration uncertainty for each method (0.04 mag for each; Rizzi et al. 2007a; Freedman 2021; Lee et al. 2024c). Separately, the blue points show a comparison between the EDD TRGB distances and the ground-based JAGB distances from Zgirski et al. (2021) using data from the InfraRed Survey Facility (IRSF), which show a somewhat reduced discrepancy of = 0.10 0.04 (stat) 0.06 (sys) mag, but with only six overlapping galaxies.
The bottom-half of Figure 1 shows four example HST CMDs used for the TRGB measurements. On each of these CMDs, we plot the measured magnitude of the TRGB (in black), and the implied magnitude of the TRGB given by the JAGB distances (in red and blue). For NGC 6822, we see excellent agreement between the two techniques. However, for the majority (3/4 cases shown, 9/13 cases total), we see that the JAGB-implied values are not consistent with the visible onset of the TRGB in these nearby galaxies. While only four examples are shown here, the remainder of the HST CMDs are available via EDD.

### v2

@@ 2 Comparison of Distance Scales
The comparison between the two samples is shown in the top-half of Figure 1. The plotted uncertainties in the JAGB distances are re-tabulated from Table 3 in Lee et al. (2024c) to include contributions from the 2MASS photometric zeropoint and 2MASS-to-FourStar correction, error on the JAGB mode, choice of smoothing scale, and a 15 uncertainty in the foreground reddening (). We also include JAGB results for WLM (Lee et al., 2021) and M33 (Lee et al., 2022), which are part of the same ground-based observing program, for a total of thirteen galaxies for comparison. The uncertainties in the TRGB distances are re-tabulated from the CMDs/TRGB catalog to include the statistical uncertainty in the tip measurement and the same 15 uncertainty in foreground reddening (). The mean difference between the two sets of distances is = 0.17 0.04 (stat) 0.06 (sys) mag. The statistical uncertainty is given by , and the systematic uncertainty is given by the quadrature sum of the underlying calibration uncertainty for each method (0.04 mag for each; Rizzi et al. 2007a; Freedman 2021; Lee et al. 2024c). Separately, the blue points show a comparison between the EDD TRGB distances and the ground-based JAGB distances from Zgirski et al. (2021) using data from the InfraRed Survey Facility (IRSF), which show a somewhat increased discrepancy of = 0.23 0.05 (stat) 0.06 (sys) mag, but with only six overlapping galaxies.
The bottom-half of Figure 1 shows four example HST CMDs used for the TRGB measurements. On each of these CMDs, we plot the measured magnitude of the TRGB (in black), and the implied magnitude of the TRGB given by the JAGB distances (in red and blue). For NGC 6822, we see excellent agreement between the TRGB measurement from Anand et al. (2021) and the JAGB value from Lee et al. (2024c). However, for the majority (3/4 cases shown, 9/13 cases total), we see that the JAGB-implied values are inconsistent with the visible onset of the TRGB in these nearby galaxies. While only four examples are shown here, the remainder of the HST CMDs are available via EDD.